OpenStudy (anonymous):

Solve the proportion. 2y / 9 = 8 / 4y

OpenStudy (anonymous):

\[2y/9 = 8/4y\] \[=> 2y * 4y = 8*9\] \[=> 8y^2 = 72\] \[=> y^2 = 72/8\] \[=> y^2 = 9\] \[=> y = \pm 3\]

OpenStudy (anonymous):

Hope you understood !!

OpenStudy (aravindg):

hope u undrstood